NetWinder Debugger Notes

Scott Bambrough, scottb@netwinder.org $Revision: 1.1.1.1 $, $Date: 1999/11/02 15:25:12 $

This document describes the debugger used on the NetWinder.

Contents

1 GNU Debugger 1 1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Known bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Insight 2 2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Known bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Miscellaneous 2 3.1 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.2 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.3 Copyright Notice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1 GNU Debugger

1.1 Introduction

The current version of GDB for the NetWinder is based on GNU release 4.18. It is a port of GDB to ARM Linux. It is still beta, but usable. I am currently in the process of submitting the patches to the GDB maintainers. Once that is done, I will concentrate on fixing the problems. The goal is to have a fully functional GDB that passes all the test suites for the release of GDB 5.0 in the new millenium.

1.2 Known bugs

Threading support is currently not functional with GLIBC’s POSIX threads implementation. GDB cannot currently load a core file. These are not really bugs, just features that have not been implemented. Tracing into shared libraries is still has some rough edges, but seems functional. I don’t have anything concrete to point at as broken, but I think it could be better. Stepping over the function prologue to point at the function code could be better. The flags are not displayed when the registers are dumped.

2 Insight

2.1 Introduction

I have also applied the changes to the Insight snapshots and have built them successfully. For those of you who don’t know Insight is GUI version of GDB, based on earlier work by Cygnus on GdbTk. For those who don’t like the command line, check it out.

2.2 Known bugs

Since Insight is an extension of GDB, all the bugs present in the ARM Linux port of GDB exist in it. Insight has not been extensively tested.

3 Miscellaneous

3.1 Author

The author and maintainer of the NetWinder Debugger Notes is Scott Bambrough (scottb@netwinder.org). Please send any comments, additions, or corrections so they can be included in the next release. The latest version of this document may be obtained from http://www.netwinder.org/~scottb/notes/

Debugger-Notes.html .

3.2 History

The first public release of this document was as an html web page.

October 21, 1999 (version 1.0): Converted web page to SGML, and updated the content.

3.3 Copyright Notice

This document is copyright (c) Scott Bambrough, 1999.

Permission is granted to make and distribute verbatim copies of this document. The copyright notice and this permission notice must be preserved on all copies.